A farmer had 15 animals in its farmyard. Some were chickens and some were cows. There were 52 legs in all. How many cows were in the farmyard?
Answer by Boreal(15235)

x=chickens
y=cows
x+y=15
2x+4y=52
-2x-2y=-30, multiplying the first equation by -2. Now add them.
2y=22
y=11 cows and 44 lege
x=4 chickens and 8 legs